Please provide a detailed response so that I may be able to understand how you came to your answer. Thank you!!!

freely trading Problem 2. (Monopolistic competition) [week 5]. Consider two countries that are in differentiated products. Each producer in the industry is subject to increasing returns to scale, derived from fixed costs of production. In particular, the cost function of a firm in the industry is linear with respect to output, with fixed costs of 100 and variable costs of 20 per unit of output. The demand function for each differentiated product is equal to
1 Q =S 1 -(P-P) 2 n where S is size of the market, P is the price charged by the producer and P is the average price in the industry. There is free entry in the industry. Assume that the size of the market is 2,000 in the Home country and 3,000 in the Foreign country HINT (in case you need it): If the demand function is: Q = A-BP, then MR = P-Q/B] a) Compute the equilibrium price that any producer will charge, as a function of the number of firms in the industry and the size of the market. b) Write down the average costs faced by any firm, as a function of the number of firms in the industry and the size of the market. c) Compute the number of firms (in the long run), the price charged for each product, and the quantity produced by each firm in the industry in the free trade equilibrium. Show it in a graph. d) Assume now that entry in the market is not free: each firm has to pay a license fee 300 to its own government, to be renewed every year, in order to participate in the market. Compute the number of firms (in the long run) in the free trade equilibrium under this new situation. e) Are consumers better off or worse off under the government licensing regime? EXPLAIN.
Answer #1

S=2000 Sp-3000 =S FC=100, VC= 20 (MC) MA = MC at Eq NOw a) O= A -BP Let A= S-SP then MR P- 81B Pa at Eqm &= SIn MR- P-28 P-DP=AC eam at &000 40=n2 6-384 n= 6 n PseignNow fC cby 300 at New m 30 New FC 400. P- a0t AC= 4cen+ 20 000 Date nA= TO 8.16 23 In foseign &-4n n2= 60/4-15 nf= N15 3.87

E) consumers are worse off , bcoz number of firms fall, so consumers get less variety of goods